OpenStudy (anonymous):

The ratio of the length of a rod and its shadow is 1: 3 . The altitude of the sun is

OpenStudy (anonymous):

my answer is 30 degree

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

i am not getting it as 30

hartnn (hartnn):

is it \(1: \sqrt 3\) ?

OpenStudy (anonymous):

just yes

hartnn (hartnn):

yeah, then the elevation angle is \(\tan^{-1}(1/\sqrt 3) = 30 ^o\)

OpenStudy (anonymous):

OKAY

OpenStudy (anonymous):

HERE ANOTHR ONE